b) What pencent of adak American males a whe than wh considered healthy c) The Cenun nepots that the pencent ofdu Amerca females with a weight higher than what is considered heaildheySeDs the of reohabilities for males and females who ane over a healdhy weighe Rules I and 2 7 Rables in Florida. Rabies is a viral de ofmited thogh t bite of a rabid animal. The virus infects the central nervous system encephalopachy and uitimately death. The Rlosda Depatmnt of Health nep the distribution of documented cases of aies foe all of 20I Species Probability 0.57 015 011 Raccoon Bar Fo Other (a) What peobability should teplacein the dismibution (b) What is the probability that a nepoted case of abies is not a accoon (c) What is the probability that a reported cane of taes is cither a hut or o Discrete probability models Examples 9.4,9.5, and 9.6 illustrate one way to assign probubilities to events A sign a probability to every individual outcome, then add these probabilities to fnd the probability of any event. This idea works well when there are only a finite (fixed and limited) number of outcomes DISCRETE PROBABILITY MODEL A probablity model with a sample space made up of a list of individal outcomes is called discrete. To assign probailities in a discrete model, list the probabilities of all the individual outcomes. These prohabalities must be numbers between O and I and must have sum 1. The probabelity of any event is the sum of the probabilities of the outcomes making up the event EXAMPLE 9.7 Hearing impairment in
Answer 9.7
Answer #1

1) Probability replaced in place "?" :

Total probability is 1 :

P(Raccoon) + P(Fat) +P( fox) +P( other) =1

0.57 + 0.15 + 0.11+ P(other) = 1

0.83 +P(other) = 1

P(other) =0.17

2) P(reported case of rabies is not raccoon) = 1 - P(reported case of rabies is raccoon)

= 1 - 0.57

= 0.43

3) P(reported case of rabies is either fat or fox) = P(fat) + P(fox)

= 0.15 + 0.11

= 0.26